Paceheco Corporation uses the weighted-average method in its process costing system. The Molding Department is the second department in its production process. The data below summarize the department’s operations in January.

Units Percent Complete with
Respect to Conversion
Beginning work in process inventory 6,050 70 %
Transferred in from the prior department during January 59,950
Completed and transferred to the next department during January 57,750
Ending work in process inventory 8,250 40 %

The accounting records indicate that the conversion cost that had been assigned to beginning work in process inventory was $35,508 and a total of $560,204 in conversion costs were incurred in the department during January. The cost per equivalent unit for conversion costs for January in the Molding Department is closest to:

  • $9.832

  • $8.587

  • $9.758

  • $9.631
Answer #1

Equivalent unit of conversion = Units completed and transferred + Units in ending inventory x completion %

= 57,750 + 8,250 x 40%

= 57,750 + 3,300

= 61,050

Total cost to account for = Beginning work in process inventory + Cost added during this month

= 35,508+560,204

= $595,712

Cost per equivalent unit for conversion = Total cost to account for / Equivalent unit of conversion

= 595,712/61,050

= $9.758

Third option is correct option.

Kindly comment if you need further assistance. Thanks‼!